# RobGC: Towards Robust Graph Condensation

### v1

@@ IV-A Experimental Settings
For a fair comparison, we follow the GCond [13] for the condensation setting. Across all GC methods compared, SGC [40] is utilised as the relay model and the downstream GNN is GCN [29]. We condense the training graph with $N$ nodes into a condensed graph with $N^{\prime}$ nodes, with the condensation ratio determined by $r=\frac{N^{\prime}}{N}$. Following GCond, the number of condensed nodes per class are selected as {5, 10, 20} for the Cora, Citeseer, and Pubmed, respectively.
@@
Hyper-parameters and Implementation. The hyper-parameters for GC methods are configured as described in their respective papers if clarified, while others are determined through grid search on the validation set. For all datasets, a 2-layer graph convolution is employed, and the width of the hidden layer is set to 256. The learning rate for the condensation process is determined through a search over the set {1e-2, 5e-3, 1e-3, 5e-4, 1e-4}. The periodic parameter $\tau$ is set as 50. For the grid search, the step parameter $c$ is adjusted to search through 10 values for the Ogbn-arxiv dataset and 20 values for all other datasets. The number of propagation for correlation matrix is set as $K=2$. The teleport probability is fixed at $\alpha=0.9$, and the iteration of label propagation is set as $l=10$. In the process of locality-based edge addition, we investigate up to 3-hop neighbors (i.e., $L$). To mitigate randomness, each experiment is repeated 5 times, and the mean accuracy and standard deviation are reported.
@@
We use the ADAM optimization algorithm to train all the models. The codes are written in Python 3.9 and the operating system is Ubuntu 16.0. We use Pytorch 1.12.1 on CUDA 12.0 to train models on GPU. All experiments are conducted on a machine with Intel(R) Xeon(R) CPUs (Gold 6326 @ 2.90GHz) and NVIDIA GeForce A40 GPUs.
@@ IV-B Denoising Performance (Q1&Q2)
We present the accuracy of various denoising methods against different levels of random noise and adversarial attacks in Table II and Table III, respectively. In these tables, Whole refers to GNNs trained on the noisy training graph and deployed on the noisy test graph. Notice that it suffers from intensive computational costs due to the large scale of the training graph. Plain represents the GNN trained on the condensed graph without denoising and deployed on the noisy test graph. In scenarios without graph noise, GC methods all maintain similar performance levels with Whole, confirming the effectiveness of GC. However, as noise levels increase, the performances drop and the gap grows between Whole and Plain, particularly notable on the dataset Ogbn-arxiv. This trend verifies that noise in the training graph significantly degrades the quality of the condensed graph produced by various GC methods.
Regarding results against random noise, the performances of GC methods are successfully improved when equipped with denoising baselines. As noise levels increase, the benefits of enhanced graph quality become more evident. Nonetheless, due to the lack of label guidance for structure optimization, these baselines inevitably introduce unreliable edges or destroy the potentially effective structures, leading to sub-optimal performance in some noise-free scenarios. Moreover, denoising baselines exhibit varying effectiveness across datasets; for instance, SVD significantly outperforms Jaccard and kNN on the Pubmed dataset across all GC methods. Our proposed RobGC consistently outperforms other baselines across different GC methods and noise levels. With the guidance of label propagation, RobGC can outperform Plain in noise-free conditions. Even at high noise levels, our proposed method maintains superior performance than other baselines by a larger margin. Remarkably, with substantial condensation ratios, RobGC achieves better results than Whole on the Citeseer when equipped with GCond. This phenomenon is attributed to the effectiveness of test-time denoising and enhanced test graph quality.
@@
TABLE III: The performance comparison of denoising methods under adversarial attack conditions. Refer to Table II for the results of Noise level=0%.
Dataset ($r$) Cora (5.80%) Citeseer (3.30%)
Noise level (%) 5 10 15 20 25 5 10 15 20 25
Whole 84.80±0.14 82.73±0.09 81.83±0.33 80.33±0.45 79.10±0.86 76.17±1.05 75.67±0.21 73.43±0.37 72.23±0.82 71.57±0.29
GCond Plain 76.83±0.33 74.80±0.65 73.00±0.29 72.23±0.17 68.17±0.26 73.30±0.36 71.57±0.33 70.60±0.22 69.27±0.42 64.40±0.43
Jaccard 77.10±0.57 76.57±0.05 76.50±0.22 74.77±0.40 72.07±0.26 73.87±0.29 71.93±0.48 72.57±0.19 68.37±0.17 67.73±0.21
SVD 77.33±0.37 75.83±0.25 74.07±0.05 72.87±0.29 70.50±0.60 74.37±0.12 73.73±0.21 71.40±0.33 71.87±0.17 69.37±0.38
kNN 77.07±0.90 75.30±0.49 74.37±0.24 72.43±0.31 71.63±1.41 74.67±0.12 72.70±0.22 71.03±0.65 69.83±0.09 67.00±0.14
Ours 78.73±0.97 77.77±0.39 76.63±0.12 75.57±0.90 73.87±0.17 75.80±0.42 75.40±0.08 75.03±0.45 74.47±0.77 73.37±0.46
GCDM Plain 72.67±0.40 71.00±0.33 68.80±0.80 64.17±0.33 62.87±0.60 73.10±0.08 71.60±0.14 68.23±0.25 67.87±0.29 62.73±0.31
Jaccard 74.67±0.94 73.07±0.62 72.20±0.28 72.17±0.54 71.27±0.63 73.57±0.12 72.47±0.19 70.67±0.25 70.50±0.16 69.73±0.41
SVD 73.67±0.31 73.07±0.47 72.47±0.60 68.40±0.24 65.17±0.25 73.43±0.05 71.90±0.16 69.80±0.28 68.87±0.19 67.60±0.08
kNN 73.03±0.33 72.17±1.66 70.53±0.05 69.93±0.21 66.90±0.67 73.70±0.73 71.87±0.29 69.67±0.37 69.13±0.33 63.17±0.52
Ours 76.67±0.49 75.43±0.05 74.77±1.01 74.60±0.08 71.63±1.16 73.90±0.08 73.40±0.22 72.97±0.19 72.83±0.05 71.83±0.05
GCEM Plain 74.23±0.38 72.97±0.05 71.53±0.52 66.90±0.22 63.10±0.29 69.47±0.41 68.43±0.05 67.67±0.33 66.77±0.26 64.43±0.90
Jaccard 76.33±0.21 74.77±0.17 72.23±0.17 70.80±0.14 69.43±0.12 70.83±0.12 70.20±0.16 69.00±0.08 68.13±0.54 66.60±0.14
SVD 75.27±0.39 73.80±0.29 72.50±0.22 69.80±0.33 69.10±0.14 71.67±0.19 70.37±0.17 69.00±0.22 68.63±0.09 66.53±0.12
kNN 75.37±0.61 73.77±0.26 72.93±0.88 69.70±1.07 68.07±0.09 70.43±0.45 70.47±0.17 69.33±0.21 68.93±0.69 67.40±0.50
Ours 77.80±0.51 75.87±1.01 73.93±0.05 71.03±0.12 70.10±0.45 73.47±0.17 72.90±0.08 72.67±0.12 72.63±0.17 71.33±0.26
Dataset ($r$) Pubmed (0.17%) Ogbn-arxiv (0.50%)
Noise level (%) 5 10 15 20 25 5 10 15 20 25
Whole 83.60±0.08 82.77±0.17 81.47±0.37 79.23±0.21 77.63±0.34 65.79±0.23 64.29±0.05 63.88±0.17 63.04±0.09 62.19±0.09
GCond Plain 77.60±0.33 75.27±0.19 74.20±0.14 72.13±0.21 69.23±0.09 54.43±0.16 51.88±0.24 46.70±0.14 45.90±0.51 45.40±0.23
Jaccard 77.40±0.24 76.07±0.05 73.70±0.24 72.40±0.50 70.37±0.19 54.76±0.44 52.27±0.20 48.25±0.63 47.95±0.24 46.39±0.36
SVD 80.13±0.29 79.87±0.33 79.50±0.43 79.27±0.26 78.33±0.29 54.52±2.92 52.57±2.38 48.74±3.36 48.51±3.74 47.82±2.97
kNN 78.03±0.17 76.67±0.26 75.50±0.36 73.73±0.37 70.80±0.37 54.62±0.81 52.94±0.87 47.94±0.87 46.38±0.73 46.02±0.91
Ours 80.57±0.29 79.90±0.29 79.77±0.26 79.50±0.13 78.73±0.37 54.83±0.56 53.11±0.34 50.76±0.65 49.88±0.21 48.23±0.43
GCDM Plain 78.13±0.96 75.40±0.41 73.90±0.93 71.93±0.41 67.93±0.38 54.45±0.28 48.03±0.95 46.69±0.18 45.37±0.72 44.47±0.32
Jaccard 78.83±0.40 76.50±0.33 74.30±0.88 72.93±0.05 69.93±0.41 54.95±0.49 50.82±0.22 50.16±0.18 48.80±0.68 48.52±0.37
SVD 78.97±0.48 78.63±0.12 77.40±0.24 76.47±0.25 76.03±1.16 54.78±0.66 48.74±0.67 47.52±1.50 46.11±1.35 46.08±1.62
kNN 78.77±0.40 76.87±0.12 75.40±0.22 72.40±0.16 69.83±0.21 54.98±0.24 49.76±0.48 48.76±0.96 47.27±0.65 47.20±0.63
Ours 79.37±0.24 79.17±0.17 78.03±0.37 77.20±0.14 77.10±0.14 55.67±0.24 53.83±0.14 52.54±0.48 50.70±0.25 50.66±0.32
GCEM Plain 75.13±1.02 64.00±0.00 63.00±0.16 62.83±1.16 61.37±0.09 55.11±0.32 52.32±0.24 51.72±0.45 50.75±0.12 47.71±0.51
Jaccard 76.07±0.19 65.23±0.09 64.83±0.17 63.30±0.99 61.87±0.12 56.21±0.52 53.10±0.17 51.93±0.44 51.38±0.33 49.71±0.16
SVD 76.67±0.26 66.77±0.05 64.13±0.12 63.53±0.19 63.13±0.58 56.30±0.88 54.13±0.17 52.40±0.28 51.53±0.41 49.53±0.49
kNN 76.83±0.29 66.17±0.09 64.83±0.56 63.80±0.57 62.07±0.34 55.66±0.58 53.85±0.64 51.95±0.56 50.95±0.96 48.77±0.55
Ours 80.23±0.21 68.60±0.29 67.20±0.33 66.33±0.41 65.13±0.05 57.80±0.21 54.66±0.48 53.67±0.43 53.00±0.33 51.92±0.34
@@ IV-C Performance Across Condensation Ratios (Q3)
Dataset $r$ Plain Jaccard SVD kNN Ours Whole
Cora 2.90% 63.73±0.73 65.60±0.24 66.97±0.59 64.20±0.14 70.33±0.12 70.90±1.13
5.80% 64.67±0.79 67.47±0.52 68.47±0.45 65.73±0.73 70.57±0.21
11.60% 65.83±0.34 67.77±0.40 69.27±0.93 66.90±0.49 72.60±0.41
Citeseer 1.65% 66.03±0.74 68.03±0.09 68.17±0.34 66.13±0.42 71.93±0.12 69.17±0.26
3.30% 66.23±0.12 68.00±0.29 68.27±0.21 67.30±0.22 72.90±0.08
6.60% 68.03±0.45 69.03±0.17 69.07±0.19 69.67±0.94 74.70±0.14
Pubmed 0.09% 68.57±0.38 72.93±0.68 74.70±0.22 72.73±0.54 76.20±0.24 77.50±0.22
0.17% 69.77±0.29 73.67±0.56 75.53±0.69 73.63±0.85 76.80±0.28
0.33% 70.83±0.31 73.70±0.22 76.40±0.08 74.37±0.25 78.93±0.12
Ogbn-arxiv 0.25% 45.10±0.51 46.82±0.16 46.03±0.55 46.97±0.21 47.93±0.37 61.24±0.08
0.50% 46.43±0.45 47.80±0.21 47.97±0.75 47.13±0.42 49.67±0.26
1.00% 46.60±0.23 48.33±0.25 48.17±0.55 47.80±0.38 51.73±0.35
TABLE V: Architecture generalizability of RobGC equipped with different graph condensation methods. Downstream GNNs are trained on condensed graphs for evaluation. The random noise is applied and noise level is 100%.
@@
Cora (5.80%) GCond+Ours 70.57±0.21 72.73±3.50 71.90±0.35 72.90±0.57
GCDM+Ours 63.77±0.60 64.33±0.09 62.30±0.62 62.40±0.08
GCEM+Ours 63.70±0.33 64.90±0.14 63.90±0.15 61.17±1.43
Citeseer (3.30%) GCond+Ours 72.90±0.08 71.13±0.71 71.57±1.81 73.40±0.41
GCDM+Ours 70.17±0.08 71.60±0.42 70.03±0.05 71.70±0.28
GCEM+Ours 68.07±0.17 69.10±0.11 68.10±0.31 69.87±0.76
Pubmed (0.17%) GCond+Ours 76.80±0.28 77.80±0.22 77.13±0.77 75.10±1.06
GCDM+Ours 77.30±0.43 75.03±0.12 75.63±0.52 76.50±0.36
GCEM+Ours 72.93±0.05 72.27±0.09 71.20±1.70 71.67±1.22
Ogbn-arxiv (0.50%) GCond+Ours 49.67±0.26 50.70±0.36 50.77±0.25 49.87±0.25
GCDM+Ours 49.27±0.36 50.73±0.29 49.17±0.18 50.04±0.21
GCEM+Ours 52.03±0.28 52.77±0.35 52.93±0.29 51.10±0.33
@@
TABLE VII: The ablation study of RobGC. GCond is deployed to condense the graph. Plain indicates the GCond result without denoising for both training and test graph. The random noise is applied and noise level is 100%.
Dataset
($r$)
@@
Plain 64.67±0.79 66.23±0.12 69.77±0.29 46.43±0.45
w/o E 68.04±0.38 71.06±0.19 74.83±0.37 48.22±0.41
w/o propagate 69.77±0.12 72.40±0.16 76.07±0.40 48.93±0.24
w/o addition 68.33±0.29 71.47±0.05 75.53±0.12 48.11±0.21
w/o deletion 68.53±0.22 71.53±0.25 76.20±0.29 47.94±0.41
Ours 70.57±0.21 72.90±0.08 76.80±0.28 49.67±0.26
IV-D Generalizability for GNN Architectures (Q4)
A critical attribute of GC is its ability to generalize across different GNN architectures, thereby allowing the condensed graph to serve as a versatile foundation for training multiple GNN models in downstream tasks. To investigate this attribute, we assess the performance of different GNN models, including GCN [29], SGC [40], GraphSAGE [20] and APPNP [41], using condensed graphs produced by GC methods combined with our proposed method. These well-trained models are then deployed on the test graph after test-time denoising and results are presented in Table V. While performance varied among the condensation methods, all GNN models tested exhibited similar levels of effectiveness within each specific GC method. This consistency underscores the inherent generalizability of GC methods. Moreover, a significant factor contributing to this success is our proposed optimization strategy based on label propagation. This strategy eliminates the introduction of additional GNN models, therefore it not only simplifies the optimization process but also ensures broad architectural generalizability, highlighting the effectiveness of our approach in maintaining high performance across various GNN architectures.
IV-E Running Time (Q5)
The denoising process of RobGC comprises four distinct steps, including: correlation matrix calculation, unreliable edge deletion, locality-based edge addition and parameter search. We evaluate the running time for each step and the total denoising and condensation processes in Table VI. Utilizing sparse tensors to store adjacency matrices enables efficient processes for indexing edges and calculating reliability. Consequently, the running time for the initial three steps is kept under 1 second. The search process investigates 10 values for each threshold on the Ogbn-arxiv dataset and 20 for all other datasets. This results in a total of 100 and 400 threshold combinations being evaluated, respectively. The table clearly shows that the search times are well-controlled across all datasets. This efficiency stems primarily from the utilization of sparse matrix multiplication during the label propagation process and the proper configuration of search steps. The comparison of total running time shows that our method achieves effective denoising results within practical time, making it a highly viable solution for real-world applications.
TABLE VIII: Comparison between statistics of noisy and optimized graphs. GCond is employed to condense the graph. The random noise is applied and the noise level is 100%.
@@ IV-E Running Time (Q5)
Noisy Optimized Noisy Optimized Noisy Optimized Noisy Optimized
#Nodes 1,028 1,028 2,208 2,208 1,827 1,827 2,827 2,827
#Edges 4,354 8,993 13,878 21,636 5,508 12,756 13,042 24,442
Sparsity (%) 0.30 0.62 0.28 0.44 0.17 0.38 0.16 0.31
Homophily 0.52 0.68 0.50 0.60 0.42 0.63 0.45 0.57
@@
Noisy Optimized Noisy Optimized Noisy Optimized Noisy Optimized
#Nodes 18,217 18,217 19,217 19,217 90,941 90,941 139,544 139,544
#Edges 151,014 200,926 168,544 219,257 1,404,384 958,136 3,143,720 2,095,861
Sparsity (%) 0.05 0.06 0.05 0.06 0.02 0.01 0.02 0.01
Homophily 0.58 0.77 0.58 0.74 0.39 0.63 0.36 0.54

### v2

@@ IV-A Experimental Settings
For a fair comparison, we follow the GCond [11] and evaluate the node classification performance of models based on classification accuracy. Across all GC methods compared, SGC [42] is utilized as the relay model and the downstream GNN is GCN [18]. We condense the training graph with $N$ nodes into a condensed graph with $N^{\prime}$ nodes, with the condensation ratio determined by $r=\frac{N^{\prime}}{N}$. Following GCond, the number of condensed nodes per class are selected as {5, 10, 20} for the Cora, Citeseer, and Pubmed, respectively.
Dataset $r$ Plain Jaccard SVD kNN ${\rm STABLE_{G}}$ Ours Whole
Cora 2.90% 63.7±0.7 65.6±0.2 67.0±0.6 64.2±0.1 67.1±0.2 70.3±0.1 70.9±1.1
5.80% 64.7±0.8 67.5±0.5 68.5±0.5 65.7±0.7 69.3±0.2 70.6±0.2
11.60% 65.8±0.3 67.8±0.4 69.3±0.9 67.5±0.5 69.5±0.4 72.6±0.4
Citeseer 1.65% 66.0±0.7 68.0±0.1 68.2±0.3 66.1±0.4 69.2±0.3 71.9±0.1 69.2±0.3
3.30% 66.2±0.1 68.0±0.3 68.3±0.2 67.3±0.2 70.0±0.3 72.9±0.1
6.60% 68.0±0.5 69.0±0.2 69.1±0.2 69.7±0.9 70.2±0.5 74.7±0.1
Pubmed 0.09% 68.6±0.4 72.9±0.7 74.7±0.2 72.7±0.5 74.9±0.4 76.2±0.2 77.5±0.2
0.17% 69.8±0.3 73.7±0.6 75.5±0.7 73.6±0.8 75.7±0.6 76.8±0.3
0.33% 70.8±0.3 73.7±0.2 76.4±0.1 74.4±0.2 75.9±0.5 78.9±0.1
Ogbn-arxiv 0.25% 45.1±0.5 46.8±0.2 46.0±0.6 47.0±0.2 47.2±0.4 47.9±0.4 61.2±0.1
0.50% 46.4±0.5 47.8±0.2 48.0±0.8 47.1±0.4 48.5±0.1 49.7±0.3
1.00% 46.6±0.2 48.3±0.2 48.2±0.6 47.8±0.4 50.1±0.5 51.7±0.4
@@
Hyper-parameters and Implementation. The hyper-parameters for GC methods are configured as described in their respective papers if clarified, while others are determined through grid search on the validation set. For all datasets, a 2-layer graph convolution is employed, and the width of the hidden layer is set to 256. The learning rate for the condensation process is determined through a search over the set {1e-2, 5e-3, 1e-3, 5e-4, 1e-4}. For our method, the periodic parameter $\tau$ is set as 50. The number of propagation for correlation matrix is set as $K=2$. The teleport probability is fixed at $\alpha=0.9$, and the iteration of label propagation is set as $l=10$. In the process of locality-based edge addition, we investigate up to 3-hop neighbors (i.e., $L$). To optimize the grid search, the step parameter $c$ is dynamically adjusted by defining the number of search candidates $h$ as:
$c=\frac{\max({\bar{\mathbb{A}}})-\min({\bar{\mathbb{A}}})}{h-1},$ (14)
where $\bar{\mathbb{A}}$ measures the reliability for each edge, as defined in Eq. (9). Consequently, the grid search is controlled by varying $h$, which is set to 10 for the Ogbn-arxiv, Flickr, Reddit, and Gamers datasets and 20 for all other datasets.
For denoising techniques, the Jaccard similarity threshold for removing dissimilar edges is chosen from {0.01, 0.02, 0.03, 0.04, 0.05}. For SVD, the reduced rank of the perturbed graph is tuned from {10, 50, 100, 200}. For kNN, the number of neighbors is selected from {1, 2, 3, 4, 5}.
@@
To mitigate randomness, each experiment is repeated 5 times, and the mean accuracy and standard deviation are reported. The codes are written in Python 3.9 and the operating system is Ubuntu 16.0. We use Pytorch 1.12.1 on CUDA 12.0 to train models on GPU. All experiments are conducted on a machine with Intel(R) Xeon(R) CPUs (Gold 6326 @ 2.90GHz) and NVIDIA GeForce A40 GPUs.
@@ IV-B Denoising Performance (Q1&Q2)
We present the accuracy of various denoising methods against different levels of random noise and adversarial attacks in Table II and Table III, respectively. In these tables, Whole Dataset refers to GNNs trained on the noisy training graph and deployed on the noisy test graph. Notice that it suffers from intensive computational costs due to the large scale of the training graph. Plain represents the GNN trained on the condensed graph without denoising and deployed on the noisy test graph. Notice that the performance of Whole Dataset and Plain differs from the results reported in the graph condensation papers [11, 27, 29] due to the distinct experimental setting of our study, which involves smaller training graphs and inductive test graphs, as shown in Table I. As noise levels increase, the performances drop and the gap grows between Whole Dataset and Plain, particularly notable on the dataset Ogbn-arxiv. This trend verifies that noise in the training graph significantly degrades the quality of the condensed graph produced by various GC methods.
Regarding results against random noise, the performances of GC methods are successfully improved when equipped with denoising baselines. As noise levels increase, the benefits of enhanced graph quality become more evident. Nonetheless, due to the lack of label guidance for structure optimization, these baselines inevitably introduce unreliable edges or destroy the potentially effective structures, leading to sub-optimal performance in some noise-free scenarios. Moreover, denoising baselines exhibit varying effectiveness across datasets; for instance, SVD significantly outperforms Jaccard and kNN on the Pubmed dataset across all GC methods. Our RobGC consistently outperforms other baselines across different GC methods and noise levels. With the guidance of label propagation, RobGC can outperform Plain in noise-free conditions. Even at high noise levels, RobGC maintains superior performance than baselines by a larger margin. Remarkably, with substantial condensation ratios, RobGC achieves better results than Whole Dataset on the Citeseer when equipped with GCond. This is attributed to the effectiveness of test-time denoising and enhanced test graph quality. Among various GC methods, GDEM achieves superior performance compared to GCond and GCDM. However, RobGC further enhances GDEM’s results, highlighting its strong competitiveness.
@@ IV-C Performance Across Condensation Ratios (Q3)
IV-D Generalizability for GNN Architectures (Q4)
A critical attribute of GC is its ability to generalize across different GNN architectures, thereby allowing the condensed graph to serve as a versatile foundation for training multiple GNN models in downstream tasks. To investigate this attribute, we assess the performance of different GNN models, including GCN [18], SGC [42], GraphSAGE [17] and APPNP [43], using condensed graphs produced by GC methods combined with our proposed method. These well-trained models are then deployed on the test graph after test-time denoising and results are presented in Table V. While performance varied among the condensation methods, all GNN models tested exhibited similar levels of effectiveness within each specific GC method. This consistency underscores the inherent generalizability of GC methods. Moreover, a significant factor contributing to this success is our proposed optimization strategy based on label propagation. This strategy eliminates the introduction of additional GNN models, therefore it not only simplifies the optimization process but also ensures broad architectural generalizability, highlighting the effectiveness of our method in maintaining high performance across various architectures.
TABLE V: Architecture generalizability of RobGC equipped with different graph condensation methods. Downstream GNNs are trained on condensed graphs for evaluation. The random noise is applied and the noise level is 100%.
@@ IV-D Generalizability for GNN Architectures (Q4)
Cora (5.80%) GCond+Ours 70.6±0.2 72.7±3.5 71.9±0.3 72.9±0.6
GCDM+Ours 63.8±0.6 64.3±0.1 62.3±0.6 62.4±0.1
GDEM+Ours 70.9±0.2 72.9±0.2 72.1±0.3 72.3±0.4
Citeseer (3.30%) GCond+Ours 72.9±0.1 71.1±0.7 71.6±1.8 73.4±0.4
GCDM+Ours 70.2±0.1 71.6±0.4 70.0±0.1 71.7±0.3
GDEM+Ours 73.2±0.2 73.3±0.4 72.3±0.5 73.6±0.5
Pubmed (0.17%) GCond+Ours 76.8±0.3 77.8±0.2 77.1±0.8 75.1±1.1
GCDM+Ours 77.3±0.4 75.0±0.1 75.6±0.5 76.5±0.4
GDEM+Ours 78.9±0.2 78.2±0.2 77.4±0.3 75.7±0.8
Ogbn-arxiv (0.50%) GCond+Ours 49.7±0.3 50.7±0.4 50.8±0.2 49.9±0.2
GCDM+Ours 49.3±0.4 50.7±0.3 49.2±0.2 50.0±0.2
GDEM+Ours 52.1±0.3 52.6±0.4 51.9±0.5 51.5±0.4
Flickr (0.5%) GCond+Ours 48.0±0.1 47.1±0.1 46.8±0.2 46.7±0.1
GCDM+Ours 47.3±0.4 46.7±0.3 46.5±0.5 46.6±0.2
GDEM+Ours 48.5±0.1 48.7±0.2 48.1±0.3 48.2±0.2
Reddit (0.1%) GCond+Ours 89.3±0.7 89.4±0.5 88.9±0.3 87.1±0.6
GCDM+Ours 89.5±0.7 88.7±0.4 89.2±0.5 89.0±0.4
GDEM+Ours 91.1±0.2 89.6±0.1 89.4±0.3 90.9±0.1
Squirrel (1.7%) GCond+Ours 24.7±1.9 25.2±1.3 23.4±2.0 23.3±1.4
GCDM+Ours 23.9±0.7 24.2±1.3 23.2±1.2 23.1±1.4
GDEM+Ours 25.5±1.9 25.2±1.5 25.5±1.2 23.6±1.6
Gamers (0.5%) GCond+Ours 57.8±0.1 54.2±0.3 55.1±0.4 58.0±0.3
GCDM+Ours 58.6±0.1 54.7±0.2 54.9±0.3 58.2±0.2
GDEM+Ours 59.5±0.1 58.1±0.3 58.8±0.2 59.1±0.2
IV-E Running Time (Q5)
The denoising process of RobGC comprises four distinct steps, including: correlation matrix calculation, unreliable edge deletion, locality-based edge addition and parameter search. We evaluate the running time for each step and the total denoising and condensation processes in Table VI. Utilizing sparse tensors to store adjacency matrices enables efficient processes for indexing edges and calculating reliability. Consequently, the running time for the initial three steps is kept under 1 second. The search process investigates 10 or 20 values for diverse datasets. This results in a total of 100 and 400 threshold combinations being evaluated, respectively. The table clearly shows that the search times are well-controlled across all datasets. This efficiency stems primarily from the utilization of sparse matrix multiplication during the label propagation process and the proper configuration of search steps. The comparison of total running time shows that our method achieves effective denoising results within practical time, making it a viable solution for real-world applications.
@@ IV-E Running Time (Q5)
TABLE VII: The ablation study of RobGC. GCond is deployed to condense the graph. Plain indicates the GCond result without denoising for both the training and test graphs. The random noise is applied and the noise level is 100%.
Dataset ($r$)
@@
Plain 64.7±0.8 66.2±0.1 69.8±0.3 46.4±0.5
w/o E 68.0±0.4 71.1±0.2 74.8±0.4 48.2±0.4
w/o propagate 69.8±0.1 72.4±0.2 76.1±0.4 48.9±0.2
w/o addition 68.3±0.3 71.5±0.1 75.5±0.1 48.1±0.3
w/o deletion 68.5±0.3 71.5±0.3 76.2±0.3 47.9±0.4
Ours 70.6±0.2 72.9±0.1 76.8±0.3 49.7±0.3
TABLE VIII: Comparison between statistics of noisy and optimized graphs. GCond is employed to condense the graph. The random noise is applied and the noise level is 100%. $\rm{Optimized_{d}}$ and $\rm{Optimized_{a}}$ denote the test graph after edge deletion and addition, respectively.
@@
Noisy Optimized Noisy $\rm{Optimized_{d}}$ $\rm{Optimized_{a}}$ Noisy Optimized Noisy $\rm{Optimized_{d}}$ $\rm{Optimized_{a}}$
#Nodes 1,028 1,028 2,208 2,208 2,208 1,827 1,827 2,827 2,827 2,827
#Edges 4,354 8,993 13,878 11,067 21,636 5,508 12,756 13,042 12,707 24,442
Sparsity (%) 0.30 0.62 0.28 0.23 0.44 0.17 0.38 0.16 0.16 0.31
Homophily 0.52 0.68 0.50 0.58 0.60 0.42 0.63 0.45 0.55 0.57
@@
Noisy Optimized Noisy $\rm{Optimized_{d}}$ $\rm{Optimized_{a}}$ Noisy Optimized Noisy $\rm{Optimized_{d}}$ $\rm{Optimized_{a}}$
#Nodes 18,217 18,217 19,217 19,217 19,217 90,941 90,941 139,544 139,544 139,544
#Edges 151,014 200,926 168,544 136,054 219,257 1,404,384 958,136 3,143,720 1,690,911 2,095,861
Sparsity (%) 0.05 0.06 0.05 0.04 0.06 0.02 0.01 0.02 0.01 0.01
Homophily 0.58 0.77 0.58 0.68 0.74 0.39 0.63 0.36 0.52 0.54